Suppose that the Gibbs Free Energy of some system is described by the following equation:

G(T,P,N)=−NkT ln ((aT^(5/2))/P)

(a) Find the entropy of this system in terms of T, P, and N.

(b) Find the relationship between V, P, and T for this system. Find one equation relating all of these variables.


(c) Determine the total energy U of this system. Express your answer as a function of T, P, and N.
Answer #1

solution :> - The Gibbs Free Energy of system is described by the following equation GIT.PN) = - NIKT in{la pt2)/03_01 (a) frv= rday I ldp OX dp aTS2 E -NKT or V = d.&-MKT 1973 arla) -p? or vs NKT — (3) - or (PV = NKT - (4)(C) we know that G = U + PV-TS oru= G-PV + TS Putting the values of opves from equations (1), ((A) & (4) si UE-NKT In { a } -